OpenStudy (anonymous):

The value of y varies directly with x2, and y = 36 when x = 3. What is the value of y when x = 5?

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

whats your direct variation setup?

OpenStudy (anonymous):

Is it y=5? im lost

OpenStudy (amistre64):

it would be useful to start at the beginning so that we have a firm grasp on how to obtian a solution. how do we set it up?

OpenStudy (anonymous):

y=x5 ?

OpenStudy (amistre64):

direct variation means that when y increases, x increases .... y = kx^2 another way of writing this is ... y/x^2 is constant for all (x,y) 36/3^2 = y/5^2 what is y?

OpenStudy (anonymous):

I got 4 by dividing 36 by 9

OpenStudy (amistre64):

good now keep going 4 = y/25

OpenStudy (anonymous):

How do i find Y?

OpenStudy (anonymous):

Would it be 5?

OpenStudy (amistre64):

is 5/25 = 4? no

OpenStudy (amistre64):

this is basic arithmetic ...

OpenStudy (amistre64):

multiply each side by ... what?

OpenStudy (anonymous):

4?

OpenStudy (amistre64):

no .... at least not effectively is 25/25 = 1?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

4 = y/25 4(25) = y(25/25) this is 1 4(25) = y*1 what can we conclude?

OpenStudy (anonymous):

That y is 100?

OpenStudy (amistre64):

y is divided into 25 parts y/25 and 4 is equal to 1/25th of y if we add up all 25 parts of y, we get 1 y ... adding up 25 parts that are each equal to 4 is just: 4(25) = 100 yes

OpenStudy (anonymous):

Is that all?

OpenStudy (amistre64):

thats all the question requires ...

OpenStudy (anonymous):

Thank you

OpenStudy (amistre64):

youre welcome